OpenStudy (anonymous):

What is the difference between power and geometric series? is it the coefficient and ratio?

OpenStudy (anonymous):

Power series: \[ \sum_{n=0}^{\infty} x^n \] Geometric series: \[ \sum_{n=0}^{\infty} \left(\frac{1}{x}\right)^n \]